# DTCLMapper: Dual Temporal Consistent Learning for Vectorized HD Map Construction

### v1

@@ Abstract
Temporal information plays a pivotal role in Bird’s-Eye-View (BEV) driving scene understanding, which can alleviate the visual information sparsity. However, the indiscriminate temporal fusion method will cause the barrier of feature redundancy when constructing vectorized High-Definition (HD) maps. In this paper, we revisit the temporal fusion of vectorized HD maps, focusing on temporal instance consistency and temporal map consistency learning. To improve the representation of instances in single-frame maps, we introduce a novel method, DTCLMapper. This approach uses a dual-stream temporal consistency learning module that combines instance embedding with geometry maps. In the instance embedding component, our approach integrates temporal Instance Consistency Learning (ICL), ensuring consistency from vector points and instance features aggregated from points. A vectorized points pre-selection module is employed to enhance the regression efficiency of vector points from each instance. Then aggregated instance features obtained from the vectorized points preselection module are grounded in contrastive learning to realize temporal consistency, where positive and negative samples are selected based on position and semantic information. The geometry mapping component introduces Map Consistency Learning (MCL) designed with self-supervised learning. The MCL enhances the generalization capability of our consistent learning approach by concentrating on the global location and distribution constraints of the instances. Extensive experiments on well-recognized benchmarks indicate that the proposed DTCLMapper achieves state-of-the-art performance in vectorized mapping tasks, reaching $61.9\%$ and $65.1\%$ mAP scores on the nuScenes and Argoverse datasets, respectively. The source code will be made publicly available at https://github.com/lynn-yu/DTCLMapper.
@@ I Introduction
High-Definition map (HD map) represents one of the cornerstones of autonomous driving tasks, which provides prior knowledge about static obstacles. The HD map is constructed in Bird’s-Eye-View (BEV) space, similar to top-view semantic understanding [1] for LiDAR data as well as camera-LiDAR fusion [2] research. Semantic maps generally use grids to depict roads on a plane, whereas vectorized maps encapsulate road structures with geometric points and lines. Although both types serve distinct purposes, vectorized maps offer a significant advantage–they require less storage while achieving higher accuracy, a particularly valuable trait for online perception tasks in autonomous driving [3].
@@
Figure 1: Difference between current temporal fusion and the proposed consistent learning solution. (a) depicts an example of the current BEV temporal fusion study [4] where features are enhanced through temporal fusion for intermediate BEV feature layers. (b) depicts the proposed DTCLMapper for BEV HD map construction. It can enrich road information from historical instance features using a consistency learning approach.
@@
The rapid advancement of BEV scene understanding research [5] has spurred significant progress in HD map construction through deep learning, ensuring accurate local maps in dynamic environments. The transformation from perspective to BEV space is one of the key points in the study of visual HD construction. Owing to the great progress of visual BEV research, fixed depth-based methods [6, 7] and random depth-based methods [8, 9] are adopted for constructing an HD map. In addition, the map decoder module is one of the core techniques for delineating the boundaries. HDMapNet [10] selects the basic residual network as the semantic and instance segmentation head. Unlike this work, where vectorized HD maps are built through postprocessing, MapTR [11] has developed an end-to-end model suitable for vectorized HD map construction. Despite their high performances, a known and essential problem in visual HD map research lies in the sparsity of information. Specifically, it is difficult to present the rich spatial information of the 3D world (or BEV space) through the 2D perspective view where the object occlusion phenomenon exists in the complex driving surroundings.
@@
Recent studies [8, 4] have shown that online temporal BEV fusion is an effective solution to tackle this challenge in the field of dynamic object detection. As shown in Fig. 1a), in these existing works, historical BEV information is aggregated to this moment through self-attention [12], or Gate Recurrent Unit (GRU) [13]. This is defined as a “temporal fusion strategy”. However, as shown in Fig. 2, we have observed that these temporal fusion strategies have little impact on improving the quality of vectorized HD maps. Specifically, the characteristics of the integrated BEV feature close to the map instance are not clearly discernible, causing the vector point to struggle to return to its precise position, thereby impacting the detection accuracy negatively. After analysis, we reckon that these methods rely on an indiscriminate fusion of temporal features. In other words, the relationships and differences among individual objects, are still not adequately explored in terms of dissimilarity and similarity.
Figure 2: Results of different temporal fusion methods. The top right sub-figure shows the comparison results of different temporal fusion methods where the proposed consistency learning solution has excellent performance for HD map construction. From the analysis of the bottom sub-figure, due to the lack of reliable features of the fused BEV features, the vector points cannot accurately regress to the correct position, resulting in a confusing vector map.
@@
To address this challenge, we move away from prior indiscriminate temporal fusion strategies and examine how the online enhancement of sparse instance features can be achieved through the consistency of instance targets across consecutive frames. This approach focuses on preserving unique characteristics while reinforcing connections between instances in successive frames. First, we present the DTCLMapper framework, which features two consistent learning components: Instance Consistent Learning (ICL) and Map Consistent Learning (MCL). These components have a progressive and complementary relationship. On the one hand, ICL is composed of a Vector Point PreSelection Module (VPPSM) and Aggregated Instance Feature Consistent Learning (AIFCL), where VPPSM provides accurate instance features for consistent learning in instances and AIFCL promotes the consistent enhancement of sparse instances. On the other hand, the MCL component enhances the model by utilizing consistent geometric positions among instances. It uses a grid map rasterized from a vectorized map to enforce consistency, with map occupancy loss as the measurement mechanism, ensuring spatial relationships are maintained and improving model consistency and generalization. Extensive experiments on public nuScenes [14] and Argoverse [15] datasets, demonstrate that our method has greatly enhanced HD map construction, setting state-of-the-art performances of $61.9\%$ and $65.1\%$ in mAP, respectively.
@@
  • A novel Instance Consistent Learning (ICL) composed of a Vector Point PreSelection Module (VPPSM) and Aggregated Instance Feature Consistent Learning (AIFCL) is proposed to improve the accuracy of map instances underlying consistent learning. Specifically, through the positive samples with high similarity and the negative samples with different labels, the consistency of complementary reinforcement learning of the same instance is promoted.
@@ II Related Work
View Transformation in BEV Understanding: In the HD map research field, the view transformation between perspective and BEV view is an important module. A variety of pipelines translating perspective views into BEV views have been proposed in the community, ranging from Inverse Perspective Mapping (IPM) [16], depth estimation [6, 7], Variational Auto-Encoder (VAE) [17] and learning-based parameter optimization methods [18, 9, 19, 8, 20]. LSS [6] estimated the depth of each pixel through the appearance relation of adjacent objects. Then the pixel features were projected into the 3D space by the intrinsic and extrinsic parameters. Naturally, the BEV feature can be obtained simply in the 3D space. Furthermore, considering the sparsity of depth estimation, BEVDepth [7] improved the accuracy of depth with the help of supervision from LiDAR information. VPN [18] investigated the potential of using Mutli-Layer Perceptron (MLP) to realize view transformation, which implicitly included depth estimation and camera parameters in the network learning. BEVFormer [8] has proposed a classic of a different kind of work, which projected 3D points directly into 2D-pixel spaces to learn BEV features. Simultaneously, this work has introduced the temporal fusion operation. Specifically, the attention operation was used to learn the historical features initially, and then the current-moment features were learned. Recently, the improvement of BEVFormer in terms of temporal fusion was presented in [21]. Instead of merging temporal before current learning, it merged temporal BEV features directly by concatenating operations on channels. Similar to this work, BEVerse [4] designed temporal fusion for BEV features, except that GRU operations are used. All these temporal works perform indiscriminate learning at the BEV feature level. Although different features can be supplemented, similar features are prone to feature redundancy. In particular, in the HD map research, it does not bring a positive impact. Unlike existing works, we propose consistent learning for temporal instances. The idea originates from the consistency of map instances, with the aim of complementing the diverse characteristics of map instances for enhancing HD map construction.
@@
Online HD Map Learning: In an HD map, road dividers, pedestrian crossing, and road boundaries are basic semantic information. In past studies, HD map was generated in an offline manner, which is time-consuming and labor-intensive work [22]. Witnessed a rapid growth of BEV understanding, HDMapNet [10] was the first to construct an online BEV HD map where MLP was leveraged to be the view transformation. It introduced the semantic HD map characterized by a grid and the vectorized HD map comprised of vector points and lines. Then, a variety of research works [23, 24, 25, 26], have made extensive efforts to enhance the performance of semantic HD map construction. Bi-Mapper [24] proposed cross-space loss assists in supervising the transformation of IPM in the camera coordinate system, alleviating the geometric distortion to be further transmitted to the global BEV space. MVMap [25] constructed a long map where an uncertainty network was used to fuse local maps. NMP [26], taking a different route, had improved the accuracy of the local semantic map with the help of a global HD map prior. In addition, OpenStreetMap, as a prior, is also widely used in map construction [27, 28]. P-MapNet [27] fused OpenStreetMap prior based on HD map prior, benefiting for building long-range road maps. The aforementioned online map research objects are divided into two categories, a semantic map and a vectorized map. Different from a semantic map, a vectorized map is composed of vector points and lines that have fewer memory requirements. Moreover, it can provide basic road information in a task about trajectory prediction [29, 30, 31]. Different from the work [10] where the vectorized HD map was constructed via postprocesses, an end-to-end vectorized map generation method is chosen for our approach.
@@
MapTR [11] leveraged Deformable DETR [36] as a decoder and can identify points with different orderings. It greatly advanced the real-time capabilities of map generation. Recently, they optimized the form of query in training, which applied a mixed strategy for each query to enhance mapping accuracy [37]. There have been many studies that have refined this work over time [38, 39, 40]. InsightMapper [38] modified the query mechanism that is translated into a hybrid generation mechanism. PivotNet [41] designed a dynamic matching to obtain key points for each vector line. InstaGraM [42] utilized graph construction to learn points and lines in a map. As sparse points lack precision in subtle places, MapVR [43] designed a rasterization-based metric to improve the performance of a vectorized map. In addition, 3D lane detection [44, 45, 46, 47] can obtain a more accurate road structure. In contrast to the above methods without temporal learning, our work explores a unique temporal processing approach for map elements, enhancing the features of element instances through temporal consistent learning.
Contrastive Learning: Contrastive learning is a type of unsupervised learning that significantly advances unsupervised tasks [48]. SimCLR [49] proposed a simple visual representation contrast learning framework. Specifically, it designed a data enhancement idea and a nonlinear contrast feature learning module, improving the quality of learning representations. However, due to SimCLR requiring a large set of negative samples in contrast learning, MOCO [50] used momentum encoders to reduce the number of negative samples. MOCOv2 [51] followed a nonlinear contrast feature learning module in SimCLR [49] to further upgrade the unsupervised learning accuracy. At the same time, contrastive learning also shines brightly in the study of supervised learning [52, 53, 54, 55]. Compared with MOCO where samples were obtained by data augmentation, these work selected samples under supervision. In these studies, contrastive learning for temporal information has also been extensively explored, especially in video instance segmentation. IDOL [54] proposed an online framework based on contrastive learning, which improved the discriminative ability of instances by associating information between temporal series. Based on it, CTVIS [55] raised an effective training strategy to learn consistency for challenging targets in consecutive frames, such as occlusion and deformation.
@@ III-A Framework
As shown in Fig. 3, DTCLMapper follows the basic framework of constructing a vectorized map, comprising an image backbone, view transformer, BEV decoder, and multi-task head. Multi-view images $I_{j}$ $(j\in[0,n])$ are encoded by a pre-trained backbone, such as a ResNet model [56]. Then the deep feature of each image is aggregated into a BEV embedding $F_{bev}\in\mathbb{R}^{H{\times}W{\times}C}$ ($H$ and $W$ are the height and width of a BEV embedding, respectively, and $C$ denotes the learning dimension) through view transformer. As this paper focuses on dual temporal learning rather than proposing a view transformer architecture, we choose the classic method as the view transformer module, BEVFormer [8] and LSS [6]. And the Deformable DETR [36] is employed for the BEV decoder. The input of decoder contains a BEV embedding, an instance query $Q\in\mathbb{R}^{N{\times}C}$($N$ is the number of instances), the initial reference points $P\in\mathbb{R}^{N{\times}N_{p}{\times}2}$ ($N_{p}$ is the fixed number of vector points in each instance) and the point query $Q_{P}\in\mathbb{R}^{N{\times}N_{p}{\times}C}$. Finally, linear layers combined with LayerNorm and ReLU are used as the task head for predicting the label and box of the instance.

### v2

@@ Abstract
Temporal information plays a pivotal role in Bird’s-Eye-View (BEV) driving scene understanding, which can alleviate the visual information sparsity. However, the indiscriminate temporal fusion method will cause the barrier of feature redundancy when constructing vectorized High-Definition (HD) maps. In this paper, we revisit the temporal fusion of vectorized HD maps, focusing on temporal instance consistency and temporal map consistency learning. To improve the representation of instances in single-frame maps, we introduce a novel method, DTCLMapper. This approach uses a dual-stream temporal consistency learning module that combines instance embedding with geometry maps. In the instance embedding component, our approach integrates temporal Instance Consistency Learning (ICL), ensuring consistency from vector points and instance features aggregated from points. A vectorized points pre-selection module is employed to enhance the regression efficiency of vector points from each instance. Then aggregated instance features obtained from the vectorized points preselection module are grounded in contrastive learning to realize temporal consistency, where positive and negative samples are selected based on position and semantic information. The geometry mapping component introduces Map Consistency Learning (MCL) designed with self-supervised learning. The MCL enhances the generalization capability of our consistent learning approach by concentrating on the global location and distribution constraints of the instances. Extensive experiments on well-recognized benchmarks indicate that the proposed DTCLMapper achieves state-of-the-art performance in vectorized mapping tasks, reaching $61.9\%$ and $65.1\%$ mAP scores on the nuScenes and Argoverse datasets, respectively. The source code is available at https://github.com/lynn-yu/DTCLMapper.
@@ I Introduction
High-Definition maps (HD maps) represent one of the cornerstones of autonomous driving tasks, which provides prior knowledge about static obstacles. The HD map is constructed in Bird’s-Eye-View (BEV) space translated from perspective image views [1], similar to top-view semantic understanding [2] for LiDAR data and camera-LiDAR fusion [3]. Semantic maps generally use grids to depict roads on a plane, whereas vectorized maps encapsulate road structures with geometric points and lines. Although both types serve distinct purposes, vectorized maps offer a significant advantage–they require less storage while achieving higher accuracy, a particularly valuable trait for online perception tasks in autonomous driving [4].
@@
Figure 1: Difference between current temporal fusion and the proposed DTCLMapper. (a) depicts an example of the current BEV temporal fusion study [5] where features are enhanced through temporal fusion for intermediate BEV feature layers. (b) depicts the proposed DTCLMapper for BEV HD map construction. It can enrich road information from historical instance features using a consistency learning approach.
@@
As a pioneering work, HDMapNet [6] employed a simple MultiLayer Perceptron (MLP) as the view transformer module and used a basic ResNet structure as the semantic and instance segmentation head. Unlike this work, where vectorized HD maps are built through postprocessing, MapTR [7] has developed an end-to-end model suitable for vectorized HD map construction. Despite their high performances, a known and essential problem in visual HD map research lies in the sparsity of information. Specifically, it is difficult to present the rich spatial information of the 3D world (or BEV space) through the 2D perspective view where the object occlusion phenomenon exists in the complex driving surroundings.
@@
Recent studies [8, 5] have shown that online temporal BEV fusion is an effective solution to tackle this challenge in the field of dynamic object detection. As shown in Fig. 1, in these existing works, historical BEV information is aggregated to this moment through self-attention [9], or Gate Recurrent Unit (GRU) [10]. This is defined as a “temporal fusion strategy”. However, as shown in the middle and bottom of Fig. 2, these temporal fusion strategies for BEV learning from scratch have little impact on improving the quality of vectorized HD maps. Specifically, the characteristics of the integrated BEV feature close to the map instance are not discernible, causing the vector point to struggle to return to its precise position, thereby impacting the detection accuracy negatively. To delve deeper into the reasons, we analyze the overlapping rate of static instances in consecutive frames, as shown in the top of Fig. 2, where the differences in short consecutive frames are relatively small. During the process of BEV learning from scratch, facing similar features, direct merging strategies make it difficult to distinguish reliable characteristics, leading to feature redundancy issues. In other words, the relationships and differences among individual objects, are still not adequately explored in terms of their dissimilarity and similarity.
Figure 2: Analysis of temporal overlapping and results of different temporal methods. Successive perspective images and coarse BEV road images through Inverse Perspective Mapping (IPM) are presented on the top left, whereas on the right are the differential results of successive frames. It can be observed that the BEV image in the short temporal range has a small difference. The middle and bottom sub-figures show comparison results of different temporal fusion methods. Intuitively, from the perspectives of point regression efficiency, map generation quality, and accuracy of three levels, the proposed temporal learning has excellent performance for HD map construction.
@@
To address this challenge, we move away from prior indiscriminate temporal fusion strategies and examine how the online enhancement of sparse instance features can be achieved through the consistency of instance targets across consecutive frames. This approach focuses on preserving unique characteristics while reinforcing connections between instances in successive frames. First, we present DTCLMapper, which features two consistent learning components: Instance Consistent Learning (ICL) and Map Consistent Learning (MCL). These components have a progressive and complementary relationship. On the one hand, ICL is composed of a Vector Point PreSelection Module (VPPSM) and Aggregated Instance Feature Consistent Learning (AIFCL), where VPPSM provides accurate instance features for consistent learning in instances and AIFCL promotes the consistent enhancement of sparse instances. On the other hand, the MCL component enhances the model by utilizing consistent geometric positions among instances. It uses a grid map rasterized from a vectorized map to enforce consistency, with map occupancy loss as the measurement mechanism, ensuring spatial relationships are maintained and improving model consistency and generalization. Extensive experiments on public nuScenes [11] and Argoverse [12] datasets, demonstrate that our method has greatly enhanced HD map construction, setting state-of-the-art performances of $61.9\%$ and $65.1\%$ in mAP, respectively.
@@
  • A novel Instance Consistent Learning (ICL) composed of a VPPSM and AIFCL is proposed to improve the accuracy of instances underlying consistent learning. Specifically, through the positive samples with high similarity and the negative samples with different labels, the consistency of complementary reinforcement learning of the same instance is promoted.
@@ II Related Work
View Transformation in BEV Understanding: In the HD map construction research field, the view transformation between perspective and BEV view is an important module. There is a wealth of excellent studies [13, 14] work in the field of perspective view for driving scene understanding. And a variety of pipelines translating perspective views into BEV views have been proposed in the community, ranging from IPM [15], depth estimation [16, 17], Variational Auto-Encoder (VAE) [18], and learning-based parameter optimization methods [19, 20, 21, 8, 22]. LSS [16] estimated the depth of each pixel through the appearance relation of adjacent objects. Then the pixel features were projected into the 3D space by the intrinsic and extrinsic parameters. Naturally, the BEV feature can be obtained simply in the 3D space. Furthermore, considering the sparsity of depth estimation, BEVDepth [17] demonstrated that supervising depth estimation with depth values from laser point clouds effectively enhances the expression of BEV features, surpassing even transformer-based frameworks. VPN [19] investigated the potential of using MLP to realize view transformation, which implicitly included depth estimation and camera parameters in the network learning. BEVFormer [8] has proposed a classic of a different kind of work, which projected 3D points directly into 2D-pixel spaces to learn BEV features. Simultaneously, this work has introduced the temporal fusion operation. Specifically, the attention operation was used to learn the historical features initially, and then the current-moment features were learned. Recently, the improvement of BEVFormer in terms of temporal fusion was presented in [23]. Instead of merging temporal before current learning, it merged temporal BEV features directly by concatenating operations on channels. Similarly, BEVerse [5] designed temporal fusion for BEV features, except that GRU operations are used. All these temporal works perform indiscriminate learning at the BEV feature level. Although different features can be supplemented, similar features are prone to feature redundancy. In particular, in the HD map research, it does not bring a positive impact. Unlike existing works, we propose consistent learning for temporal instances. The idea originates from the consistency of map instances, with the aim of complementing the diverse characteristics of map instances for enhancing HD map construction.
@@
Online HD Map Learning: In an HD map, road dividers, pedestrian crossing, and road boundaries are basic semantic information. In past studies, HD map was generated in an offline manner, which is time-consuming and labor-intensive work [24]. Witnessed a rapid growth of BEV understanding, HDMapNet [6] was the first to construct an online BEV HD map where MLP was leveraged to be the view transformation. It introduced the semantic HD map characterized by a grid and the vectorized HD map comprised of vector points and lines. Then, a variety of research works [25, 26, 27, 28, 29], have made extensive efforts to enhance the performance of semantic HD map construction. The aforementioned online map research objects are divided into two categories, a semantic map and a vectorized map. Different from a semantic map, a vectorized map is composed of vector points and lines that have fewer memory requirements. Moreover, it can provide basic road information for tasks such as trajectory prediction [30, 31, 32] and navigation [33] of intelligent vehicles. Unlike the work [6], where the vectorized HD map was constructed through post-processing, the proposed method employs an end-to-end vectorized map generation method.
@@
MapTR [7] leveraged Deformable DETR [38] as a decoder and can identify points with different orderings. It greatly advanced the real-time capabilities of map generation. Recently, they optimized the form of query in training, which applied a mixed strategy for each query to enhance mapping accuracy [39]. There have been many studies that have refined this work over time [40, 41, 42]. InsightMapper [40] modified the query mechanism that is translated into a hybrid generation mechanism. PivotNet [43] designed a dynamic matching to obtain key points for each vector line. InstaGraM [44] utilized graph construction to learn points and lines in a map. As sparse points lack precision in subtle places, MapVR [45] designed a rasterization-based metric to improve the performance of a vectorized map. In addition, 3D lane detection [46, 47, 48, 49] can obtain a more accurate road structure. Recently, some scholars have explored the fusion and tracking of large-scale temporal memory buffers. StreamMapNet [50] introduced a streaming strategy, which is different from our stacking strategy. It blends all the visible features of the past at the feature level and joins high-confidence instances from historical temporal data. MapTracker [51] supervised the association of map instances and fused traceable objects in historical frames from BEV feature warming, instance vector warming, and joint training perspectives. In contrast to the two methods, we focus on short-term BEV features learned from scratch. Our work introduces a unique temporal processing approach during the training phase, enhancing the features of element instances through temporal consistent learning and bypassing complex temporal parameter calculations in the testing phase, which is more flexible and suitable for real-world online HD mapping.
Contrastive Learning: Contrastive learning is a type of unsupervised learning that significantly advances unsupervised tasks [52]. SimCLR [53] designed a data enhancement idea and a nonlinear contrast feature learning module, improving the quality of learning representations. However, due to SimCLR requiring a large set of negative samples in contrast learning, MOCO [54] used momentum encoders to reduce the number of negative samples. MOCOv2 [55] followed a nonlinear contrast feature learning module in SimCLR [53] to further upgrade the unsupervised learning accuracy. At the same time, contrastive learning also shines brightly in the study of supervised learning [56, 57, 58, 59]. Compared with MOCO where samples were obtained by data augmentation, these work selected samples under supervision. In these studies, contrastive learning for temporal information has also been extensively explored, especially in video instance segmentation. IDOL [58] proposed an online framework based on contrastive learning, which improved the discriminative ability of instances by associating information between temporal series. Based on it, CTVIS [59] raised an effective training strategy to learn consistency for challenging targets in consecutive frames, such as occlusion and deformation.
@@ III-A Framework
As shown in Fig. 3, DTCLMapper follows the basic framework of constructing a vectorized map, comprising an image backbone, view transformer, BEV decoder, and multi-task head. Multi-view images $I_{j}$ $(j\in[0,n])$ are encoded by a pre-trained backbone, such as ResNet (Res) [60]. Then the deep feature of each image is aggregated into a BEV embedding $F_{bev}=Prob(I_{j},VT)$ $(F_{bev}\in\mathbb{R}^{H{\times}W{\times}C})$ ($H$ and $W$ are the height and width of a BEV embedding, respectively, and $C$ denotes the learning dimension) through View Transformer (VT). As this paper focuses on dual temporal learning rather than proposing a view transformer architecture, we choose the classic method as the view transformer module, BEVFormer ($Prob_{B}$[8] and LSS ($Prob_{L}$[16].
$Z_{c}*\left[\begin{array}[]{l}u\\ v\\ 1\\ \end{array}\right]=T_{in}\cdot\left[\begin{array}[]{l}X_{c}\\ Y_{c}\\ Z_{c}\\ \end{array}\right]\left[\begin{array}[]{l}X_{c}\\ Y_{c}\\ Z_{c}\\ 1\\ \end{array}\right]=T_{ex}\cdot\left[\begin{array}[]{l}X_{w}\\ Y_{w}\\ Z_{w}\\ 1\\ \end{array}\right],\vspace{-0.15cm}$ (1)
$Prob_{B}=\sum_{i=0}^{8}F_{A}(Res(I_{j}),P3(T_{in},T_{ex},[X_{w},Y_{w},Z_{i}]^{\rm T})),$ (2)
$Prob_{L}=F_{B}(Res(I_{j}),P2(T_{in},T_{ex},[u,v]^{\rm T},{Z_{c}^{p}})),$ (3)
where $T_{in}$ and $T_{ex}$ are intrinsic and extrinsic matrixes of each camera. $X_{w}$, $Y_{w}$ and $Z_{i}$ are in the ego coordinate system. $X_{c}$, $Y_{c}$, and $Z_{c}$ are in the camera coordinate system. $u$ and $v$ are the pixel coordinate system. ${Z_{c}^{p}}$ is the value of predicted depth. $P3$ is VT from 3D points to 2D images. $F_{A}$ is deformable cross-attention learning at 2D features to obtain corresponding 3D features. $P2$ is VT from 2D pixels to 3D points. $F_{B}$ is the projection of 2D features to corresponding 3D points based on ${Z_{c}^{p}}$. And the Deformable DETR [38] is employed for the BEV decoder. The input of decoder contains a BEV embedding, an instance query $Q\in\mathbb{R}^{N{\times}C}$($N$ is the number of instances), the initial reference points $P\in\mathbb{R}^{N{\times}N_{p}{\times}2}$ ($N_{p}$ is the fixed number of vector points in each instance) and the point query $Q_{P}\in\mathbb{R}^{N{\times}N_{p}{\times}C}$. Finally, linear layers combined with LayerNorm and ReLU are used as the task head for predicting the label and box of the instance $Ins_{i}$.